OpenStudy (vannah):

what is an equation of the line that passes through the pointss(3,-3) and has a slope of 2????

OpenStudy (bahrom7893):

y+3 = 2(x-3)

OpenStudy (bahrom7893):

y = 2x-6-3 y = 2x-9

OpenStudy (aroub):

y=mx+b m=2 (3,-3) -3=2(3)+b -3=6+b -6-3=b -9=b y=2x-9

OpenStudy (vannah):

thnkss